OpenStudy (anonymous):

Choose the product. a 2 b 3 · 6ab 2 6a 2b 5 6a 2b 6 6a 3b 5

geerky42 (geerky42):

Is that \(\Large a^2\cdot b^3\cdot 6ab^2\)?

OpenStudy (anonymous):

yes

geerky42 (geerky42):

Use product of power property. You just add exponents together For example: \(\Large c^ac^b = c^{a+b}\)

OpenStudy (anonymous):

Can I just get the answer, I'm mathematically illiterate and need to pass summer school.

geerky42 (geerky42):

I'm sure you can do addition. You just add exponents together

geerky42 (geerky42):

So what \(a^2\cdot a\) would be?

geerky42 (geerky42):

\(a^2\cdot a\) is same as \(a^2\cdot a^1\)

geerky42 (geerky42):

So you just add 1 and 2 together. Do same to b.